ASSIGNMENT

1. Determine the capitalized cost of an equipment costing P2M with and annual maintenance of P200,000.00 if money is worth 20% per annum.

2. A dam will have a first cost of $5,000,000 an annual maintenance cost of $25,000 and minor reconstruction costs of $100,000 every five years. At an interest rate of 8% per year, the capitalized cost of the dam is?

3. A P100,000, 6% bond, pays dividend semi-annually and will be redeemed at 110% on July 1, 1999. Find its price if bought on July 1, 1996, to yield an investor 4%, compounded annually.

4. Determine the annual cost of a structure that requires P15 million to build with a salvage value of P2 million after 12 years if the interest rate is 6% compounded annually.

5. At 6%, find the capitalized cost of a bridge whose cost is P250M and life is 20 years, if the bridge must be partially rebuilt at a cost of P100M at the end of each 20 years.

Guys assignement.. please don`t answer if you cant answer all...
Answer #1

1.

I = 20% = 0.2

Capitalized cost = -2000000 - 200000 / 0.2

= -2000000 - 1000000

=-3000000

2. Capitalized cost = -5000000 -25000 / 0.08 - 100000 * (A/F, 8%,5) / 0.08

= -5000000 -25000 / 0.08 - 100000 * 0.170456 / 0.08

= - 5525570.57

Pls upload rest of the questions separately as per Chegg policy